OpenStudy (anonymous):

find solutions 2sinxcosx+cosx=0

OpenStudy (anonymous):

Start by factoring the equation \[\cos(x) (2\sin(x)+1) = 0\]

OpenStudy (anonymous):

Since multiplying by 0 will give 0, set both of them to equal 0. \[\cos(x) = 0\]\[2\sin(x)+1 = 0\]

OpenStudy (anonymous):

After simplifying the second equation you should get\[\sin(x)=-\frac{ 1 }{ 2 }\]Now solve for x for both equations